A survey of 100 similar-sized hospitals revealed a mean daily census in


the pediatrics service of 27


with a standard deviation of 6.5. Do these data provide sufficient evidence to


indicate that the


1. population mean is greater than 25? Let alpha = 0.5

1
Expert's answer
2022-03-24T05:49:57-0400

"H_0:\mu \leqslant 25\\H_1:\mu >25\\T=\sqrt{n}\frac{\bar{x}-\mu}{s}=\sqrt{100}\frac{27-25}{6.5}=3.07692\sim t_{n-1}=t_{99}\\P-value:\\P\left( T>3.07692 \right) =F_{t,99}\left( -3.07692 \right) =0.00135"

Since the P-value is less than the significance level, the population mean is greater than 25.